OpenStudy (anonymous):

word problem: Jenna's ipod has 250 songs on it and she adds 3 new songs every week. Amy's ipod has 320 songs on it and she adds 1 new song every week. In how many weeks will Jenna and Amy have the same amount of songs on their iPods

OpenStudy (anonymous):

how do you solve after 250+3x=320+x

OpenStudy (anonymous):

After \(n\) weeks, how many songs will each of them have? What is the relationship between those two expressions when they have the same amount of songs?

OpenStudy (anonymous):

I'm still pretty confused..

OpenStudy (anonymous):

To solve \(250+3x=320+x\), start by subtracting \(x\) and then \(250\) from both sides